OpenStudy (anonymous):

how do you solve e^(2x)+2e^(x)-24=0

OpenStudy (slaaibak):

hint: set k = e^x now it's a quadratic. solve for k, and set equal to e^x to find x.

OpenStudy (anonymous):

what do you mean?

OpenStudy (raden):

factor out that, gives e^(2x)+2e^(x)-24=0 (e^x-4)(e^x+6)=0 for zeroes, take each factors and set be zero. so e^x - 4 = 0 or e^x + 6 = 0 solve for x ....

OpenStudy (anonymous):

the 24 isnt in the exponent though....

OpenStudy (raden):

no, 24 just a constant